If Kp = 307.7 for the equilibrium reaction below at 326.2 °C, find Ko CH_(9) +202(g) = CO2(g) +2H20(1)
Answer #1

T= 326.2 oC
= (326.2+273) K
= 599.2 K
Δ n = number of gaseous molecule in product - number of gaseous molecule in reactant
Δ n = -2

use:
Kp= Kc (RT)^Δn
3.077*10^2 = Kc *(0.08206*599.2)^(-2)
Kc = 7.439*10^5
Answer: 7.44*10^5